# GroupNL: Low-Resource and Robust CNN Design over Cloud and Device

### v1

@@ 3.4 Resource and Training Acceleration Analysis
[…]
Noticeably, the relation is $\mathcal{G}_{\mathrm{groupnl}}^{\mathrm{sparse}}\leq\mathcal{G}_{\mathrm{groupnl}}=c_{\mathrm{in}}c_{\mathrm{out}}k^{2}/r<\mathrm{min}\{{\mathcal{G}}_{\mathrm{cnn}},{\mathcal{G}}_{\mathrm{ghostnet}},{\mathcal{G}}_{\mathrm{sinefm}}\}$, i.e., the number of gradients for optimization in GroupNL is minimal compared to the rest CNN design methods of ${\mathcal{G}}_{\mathrm{cnn}}$, ${\mathcal{G}}_{\mathrm{ghostnet}}$, and ${\mathcal{G}}_{\mathrm{sinefm}}$. The sparse GroupNL also significantly reduces the gradient exchange by a factor of $\frac{1}{r\cdot\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{out}}/r)}$ compared to the ${\mathcal{G}}_{\mathrm{cnn}}$. As the number of GPUs scales, this scatter/reduction gap compared to vanilla CNNs also grows larger thus accelerating the training.
It is notable that, it seems that GhostNet and SineFM also have a small number of gradients with ${{\mathcal{G}}_{\mathrm{ghostnet}}<\mathcal{G}}_{\mathrm{cnn}}$ and ${\mathcal{G}}_{\mathrm{sinefm}}<{\mathcal{G}}_{\mathrm{cnn}}$, can also accelerate training in theory. However, the evaluation in Table I shows that the GhostNet and SineFM are even slower than the vanilla CNN. For this phenomenon, we consider this to be due to the introduction of more computational nodes involved in the backward updating, such as DWConv and Norm, which leads to slower training instead. Our proposed GroupNL, on the other hand, does not introduce new computational nodes through in-place operations such as $\tt{torch.cat}$ and $\tt{torch.split}$, and copy operations such as $\tt{torch.repeat}$ and $\tt{torch.repeat\_interleave}$, thus ensures acceleration of training.
3.5 Discussion
(ii) More training efficient. GroupNL generates the number of feature maps without extra computational nodes like DWConv and Norm, effectively reducing the number of gradients exchanging in the multi-GPUs training.
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets
We evaluate the effectiveness of the proposed method using five widely used datasets. CIFAR-10, GTSRB, and ImageNet-1K are standard datasets. CIFAR-10-C and Icons-50 are used as corrupted datasets. Among them, noise is introduced in the CIFAR-10-C dataset, and the training data and test data styles of Icons-50 are quite different.
GTSRB [36] is a German traffic sign dataset for traffic sign recognition. It comprises 39,209 training images and 12,630 test images from 43 categories. The images in GTSRB have varying light conditions and rich backgrounds. We set the resolution of all images to $32\times 32$ pixels.
ImageNet [38] is a large-scale classification dataset to verify computer vision performance. ImageNet-1K is a subset of the ILSVRC ImageNet, which contains 1.28 billion training images and 50K validation images from 1K different classes. The resolution of each image is $224\times 224$ pixels.
4.1.2 Baseline CNNs
We compare the following nonlinear transformation and feature maps generated-based methods in CNNs.
MonoCNN [5]: In MonoCNN, the convolutional layer is specified with only a few seed filter parameters updated during model training. The remaining filters in this layer are generated based on the seed filters and monomial function. The hyperparameters of the monomial function are randomly generated and non-learnable in training.
GhostNet [18]: In GhostNet, a few filters are first used to generate some feature maps. Then, utilizing $d\times d$ DWConv operations to generate the remaining feature maps based on these generated feature maps.
SineFM [19]: In the SineFM-based CNN model, a small number of seed convolutional filters are first specified in the convolutional layer. Then, seed feature maps are generated based on the seed filters and input data. Subsequently, a set of expanding feature maps is generated based on the seed feature map, $t$-sinusoidal functions, and a $t$-sets of BatchNorm. Finally, utilizing the $1\times 1$ DWConv to align the number of expanding features to the generated features.
To ensure a comprehensive comparison, all baseline methods are based on the following CNN models: ResNet-18/-34/-101 [2] and VGG11 [1]. To make the effect of training acceleration more clear, we evaluate the training speed of the larger model, i.e., the ResNet-101.
4.1.3 Implementation Details
Training Platforms and Libraries: The experiments are performed on the servers with NVIDIA RTX 2080Ti 11GB and RTX 4090 GPUs 24GB. We implement the models in PyTorch 1.12 for the NVIDIA platforms with CUDA 11.1.
Hyperparameters Settings of Baseline CNNs: For the hyperparameters setting, we follow the suggestion from the original papers [5, 18, 19]. The specific description of model hyperparameters is shown in Table 5.
(i) MonoCNN [5] uses the monomial function as a filter generating function and samples a random and continuous exponent $\eta$ between $[1,7]$, and the exp_factor is $\{2,4,8,16\}$;
(ii) GhostNet [18] uses $3\times 3$ grouped convolution with $d=3$ as the cheap operation for generating the feature maps and sets the reduction ratio to 2;
(iii) SineFM [19] uses a sinusoidal function as the feature map generator function with random hyperparameters, the real numbers of period $\omega$ and shift $\phi$, which ranges are between $[1,2]$ and $[1,5]$, respectively. The number of the NLFs with different hyperparameter settings in SineFM is $t=5$, which means five times expanding the seed features as described in Eq. 7.
Experimental Settings: For the settings of training in CIFAR-10/GTSRB/Icons-50, the learning rate is 0.1, and the batch size is 128, with the SGD optimizer and cosine annealing scheduler. For the training in ImageNet-1K, the learning rate is 0.001, and the batch size is 1024, with the Adam optimizer and cosine annealing scheduler.
In addition, the training evaluation is compared for the larger model, i.e., ResNet-101 in different parallel technologies in PyTorch library API [39], including the Data Parallelism (DP) with $\tt{torch.nn.DataParallel}$, the Distributed Data Parallelism (DDP) [40] with $\tt{torch.nn.parallel.DistributedDataParallel}$, and the Automatic Mixed Precision (AMP) [41] with $\tt{torch.cuda.amp}$. We also follow the linear scaling rule [42] to scale the number of parallelism in training, which is scaling the learning rate by the batch size. In addition, we also set up the same scale of seed feature map numbers in GhostNet, SineFM, and GroupNL with $r=2$.
TABLE 5: Hyperparameter setting of CNNs methods.
Method Hyperparameters Settings
Reduction Coefficient & Random Range
MonoCNN $\mathrm{exp\_factor}=\{2,4,8,16\}$ $\eta\in[1,7]$
GhostNet $r=2$ $d=3$
SineFM $r=2$ $t=5,\quad\omega\in[1,2],\quad\phi\in[1,5]$
GroupNL $r=2$ $g=4,\quad\omega\in[1,2],\quad\phi\in[1,5]$
4.2 Experimental Results
4.2.1 Reduction of Network Traffic and FLOPs
Figure 5: Number of learnable parameters for different models with different methods.
Figure 6: FLOPs for different models with different methods.
Since the non-learnable hyperparameters of NLFs are randomly initialized and remain unchanged, these parameters can be saved and reproduced with a random number. Therefore, the number of learnable parameters in the CNN model can be approximated to the number of model parameters that need to be transmitted when deploying the model on cloud-assisted IoT devices.
Fig. 5 and Fig. 6 show the number of learnable parameters and FLOPs of different CNN models for different methods. Without using sparse technology, the number of learnable parameters of GhostNet, SineFM, and GroupNL is similar, which is much less than the number of learnable parameters of vanilla CNN. As shown in Fig. 5, When based on VGG11, ResNet-18, and ResNet-34 architectures, GhostNet, SineFM, and GroupNL have a similar number of learnable parameters. When based on ResNet-101 architecture, GroupNL has the least amount of learnable parameters due to the user of sparse technology. As shown in Fig. 6, in terms of FLOPs, there are similar findings as in the amount of learnable parameters.
4.2.2 Robustness
This section evaluates the robustness of the GroupNL-based CNN on three types of datasets, including standard datasets (i.e., CIFAR-10 and GTSRB), the corrupted dataset (i.e., CIFAR-10-C), and data under different styles (i.e., Icons50).
1) Standard Data.
TABLE 6: Comparison of image classification results on CIFAR-10 and GTSRB datasets.
Methods CIFAR-10 Top-1 Acc (%) GTSRB Top-1 Acc (%)
ResNet-18 ResNet-34 ResNet-101 VGG11 ResNet-18
Vanilla CNN 95.04 95.24 95.48 91.89 97.89
MonoCNN 92.05 92.26 94.97 88.84 97.66
GhostNet 93.36 93.90 95.04 92.65 95.76
SineFM 94.36 94.77 95.06 91.27 98.18
GroupNL 95.12 95.45 95.51 91.62 98.31
Table 6 shows the comparison of image classification results of different CNN models on CIFAR-10 and GTSRB datasets under different methods. We observe that:
(i) The accuracy of the GroupNL-based CNN exceeds that of all alternatives to the standard CNN model. As shown, when using ResNet-34 as the backbone, on the CIFAR-10 dataset, GroupNL’s accuracy is 0.68% higher than SineFM’s, 1.55% higher than GhostNet’s, and 3.19% higher than MonoCNN’s. When using VGG11 as the backbone, on the CIFAR-10 dataset, GroupNL’s accuracy is 0.35% higher than SineFM’s. The main reason is that GroupNL generates diversified feature maps using the NLFs with different hyperparameter configurations. In contrast, SineFM generates feature maps using a single NLF and is unable to generate diverse feature maps. GhostNet first generates some feature maps using standard convolution operations and then uses cheap linear operations to generate other feature maps based on these generated feature maps, which limits the diversity of the feature maps. Benefiting from its ability to generate diverse feature maps, GroupNL achieves the highest accuracy among many alternatives.
(ii) The GroupNL-based CNN models outperform the vanilla CNN models in most cases. As shown in Table 6, GroupNL performs slightly worse than the vanilla CNN only when using VGG11 as the backbone. When using ResNet-18/-34/-101, the accuracy of GroupNL exceeds that of the vanilla CNN. This is mainly due to the ability of GroupNL to generate diverse feature maps.
2) Corrupted Data.
TABLE 7: Robustness to common observable corruptions. Our results are highlighted with shading.
Method Noise Blur Weather Digital mean
Vanilla CNN $52.60$ $72.78$ $84.55$ $80.70$ $72.58$
MonoCNN $56.36$ $63.24$ $77.79$ $75.17$ $68.14$
GhostNet $59.26$ $72.57$ $82.99$ $79.95$ $73.69$
SineFM $53.77$ $68.60$ $84.63$ $80.94$ $71.99$
GroupNL $57.55$ $72.96$ $85.25$ $80.68$ $74.11$
Method Noise Blur Weather Digital mean
Vanilla CNN $53.72$ $71.70$ $85.08$ $82.27$ $73.19$
MonoCNN $57.16$ $63.50$ $78.78$ $75.81$ $68.81$
GhostNet $60.30$ $74.57$ $84.10$ $80.71$ $74.92$
SineFM $59.20$ $71.72$ $86.11$ $81.83$ $74.72$
GroupNL $59.84$ $73.64$ $86.01$ $81.26$ $75.19$
Method Noise Blur Weather Digital mean
Vanilla CNN $53.97$ $73.05$ $86.02$ $81.90$ $74.78$
MonoCNN $49.03$ $74.36$ $86.60$ $82.00$ $74.26$
GhostNet $55.03$ $74.00$ $86.13$ $82.45$ $74.40$
SineFM $58.33$ $72.08$ $85.08$ $81.44$ $74.23$
GroupNL $57.90$ $71.22$ $86.45$ $81.17$ $75.04$
Method Noise Blur Weather Digital mean
Vanilla CNN $65.37$ $75.45$ $82.06$ $80.48$ $75.84$
MonoCNN $42.74$ $68.73$ $78.08$ $76.58$ $66.53$
GhostNet $52.91$ $70.28$ $80.05$ $78.65$ $70.47$
SineFM $62.31$ $72.46$ $82.46$ $82.61$ $74.96$
GroupNL $64.10$ $77.24$ $83.46$ $83.87$ $77.17$
Table 7 shows the performance comparison of different models for each method on the corrupted data. We observe that:
The GroupNL-based CNN models achieve the highest accuracy. As shown in Table 7, when ResNet-34 is used as the backbone, the accuracy of GroupNL is 0.47% higher than that of SineFM, 6.38% higher than that of MonoCNN, 0.27% higher than that of GhostNet, and even 2% higher than the accuracy of standard ResNet-34. When VGG11 is used as the backbone, the accuracy of GroupNL is 2.21% higher than that of SineFM, 10.64% higher than that of MonoCNN, 6.7% higher than that of GhostNet, and 1.33% higher than the accuracy of standard VGG11. There are three reasons for this: (i) GroupNL sets some model parameters to be non-learnable, so that the model parameters are not completely obtained through training by inputting training data, thus making the model have stronger generalization ability. (ii) The rules of the NLF can regularize the model and improve the generalization ability of the model. (iii) NLFs equipped with different hyperparameters can generate diverse feature maps and improve the knowledge expression ability of the model. Therefore, GroupNL achieves the highest accuracy.
The model’s ability to generate diverse feature maps has a great impact on the robustness of the model. As shown, with ResNet-18 as the backbone, GroupNL’s accuracy is 2.12% higher than SineFM’s. With ResNet-34 as the backbone, GroupNL’s accuracy is 0.47% higher than SineFM’s. With ResNet-101 as the backbone, GroupNL’s accuracy is 0.81% higher than SineFM’s. With the same number of parameters and computational complexity, the difference between GroupNL and SineFM is that the former can generate a variety of feature maps. This shows that increasing the diversity of feature maps can effectively improve the robustness of the CNN model.
TABLE 8: Robustness to different styles. Our results are highlighted with shading.
Method Apple Facebook Google Samsung mean
Vanilla CNN $93.66$ $87.64$ $83.20$ $80.28$ $86.19$
MonoCNN $93.02$ $85.37$ $81.60$ $78.59$ $84.65$
GhostNet $92.07$ $85.09$ $78.98$ $77.61$ $83.43$
SineFM $94.84$ $88.72$ $84.96$ $82.27$ $87.70$
GroupNL $94.52$ $91.05$ $86.27$ $84.36$ $89.05$
Method Apple Facebook Google Samsung mean
Vanilla CNN $92.77$ $86.26$ $83.81$ $82.44$ $86.32$
MonoCNN $91.46$ $85.88$ $81.55$ $79.23$ $84.53$
GhostNet $89.87$ $84.70$ $79.61$ $73.76$ $81.98$
SineFM $94.74$ $88.72$ $84.88$ $81.46$ $87.45$
GroupNL $95.19$ $90.98$ $85.93$ $84.36$ $89.12$
Method Apple Facebook Google Samsung mean
Vanilla CNN $93.64$ $87.70$ $83.64$ $83.13$ $87.03$
MonoCNN $94.84$ $87.64$ $83.62$ $82.71$ $87.20$
GhostNet $93.50$ $88.53$ $83.62$ $84.50$ $87.54$
SineFM $90.25$ $83.65$ $79.69$ $74.67$ $82.06$
GroupNL $93.98$ $90.15$ $85.51$ $83.89$ $88.38$
Method Apple Facebook Google Samsung mean
Vanilla CNN $92.54$ $86.68$ $83.67$ $80.34$ $85.81$
MonoCNN $89.55$ $81.10$ $79.82$ $75.41$ $81.47$
GhostNet $91.40$ $88.15$ $80.71$ $79.33$ $84.90$
SineFM $94.74$ $89.87$ $85.59$ $82.17$ $88.09$
GroupNL $94.23$ $91.05$ $85.30$ $82.13$ $88.18$
3) Data under different styles: We obtain conclusions similar to those for the corrupted data; the GroupNL-based CNN models achieve the highest accuracy under all backbones. As shown in Table 8, when ResNet-18 is used as the backbone, the accuracy of GroupNL is 1.35% higher than that of SineFM, 4.4% higher than that of MonoCNN, 5.62% higher than that of GhostNet, and even 2.86% higher than the accuracy of standard ResNet-18. when ResNet-101 is used as the backbone, the accuracy of GroupNL is 6.3% higher than that of SineFM, 1.18% higher than that of MonoCNN, 0.84% higher than that of GhostNet, and even 1.35% higher than the accuracy of standard ResNet-18. The reason for this is the same as that of our corrupted data; that is, (i) GroupNL makes the model parameters not completely updated by the input training data; (ii) the nonlinear transformation function in GroupNL regularizes the model; (iii) GroupNL can generate diverse feature maps. Therefore, GroupNL has the highest accuracy.
4.2.3 Training Acceleration
TABLE 9: Comparison of CIFAR-10 Training Time in Different CNN methods in RTX 2080Ti GPUs.
Tech. Parallelism No. of GPUs LR BS Method Top-1 Acc (%) Training Time Accelerated Ratio
Data Parallel $\times$1 2-GPUs 0.1 64$\cdot$2 Vanilla CNN 95.48 9.3 Hours -
MonoCNN 94.97 9.2 Hours +0%
GhostNet 95.04 12.7 Hours -37%
SineFM 95.06 15.4 Hours -66%
GroupNL 95.51 8.3 Hours +11%
(i) Compared to vanilla ResNet-101, GroupNL-based ResNet-101 achieves 11% speedup while achieving the higher accuracy, as shown in Table 9. This is because the GroupNL method uses the NLFs to generate feature maps, reducing the number of operations based on sliding windows to generate feature maps. The FLOPs of generating feature maps based on nonlinear transformation are significantly less than those of generating feature maps based on sliding windows (as shown in Fig. 6). Moreover, since only two GPUs are used, the GroupNL mainly accelerates model training by reducing model calculation time.
(ii) Although the GhostNet and SineFM have fewer parameters and computational complexity than vanilla CNN, the training time is longer, as shown in Table 9. As we analyzed in Section 2.1, although cheap linear transformations and normalizations require less computation, they are slow to compute and optimize. Incorporating these operations into the model will significantly slow down the model’s training speed.
We then further compare the speedup of GroupNL and vanilla CNN with different numbers of GPUs under Data Parallel (DP) and Distributed Data Parallel (DDP).
1) Training Evaluation in DP. For the multi-GPUs evaluation using DP technology, Table 10 and Fig. 7 show the experimental settings and training results, respectively. We observe that:
TABLE 10: Training Settings in DP Evaluation.
Platform Parallelism No. of GPUs LR No. of BS
RTX 2080Ti baseline 1-GPU 0.1 128$\cdot$1
$\times$1 2-GPUs 0.1 64$\cdot$2
$\times$2 2-GPUs 0.2 128$\cdot$2
$\times$3 3-GPUs 0.3 128$\cdot$3
$\times$4 4-GPUs 0.4 128$\cdot$4
With more GPUs, the GroupNL has a larger training speed-up ratio. As shown in Fig. 7, when the number of GPUs is 4, GroupNL has the largest speed-up ratio compared to the vanilla CNN. This is because when the number of GPUs is large in DP, the parameter transmission time between multiple GPUs will increase significantly when training vanilla CNN. In contrast, the non-trainable parameters in the GroupNL do not need to be transmitted and are less affected by the parameter transmission between multiple GPUs. Therefore, when a large number of GPUs are used, GroupNL has a higher training speedup ratio.
2) Training Evaluation in DDP. For the multi-GPUs evaluation with DDP technology, Table 11 and Fig. 8 show the experimental settings and evaluated results of DDP training with AMP, respectively. We observe that: Since DDP uses Ring-AllReduce to reduce parameter transmission between multiple GPUs, the GroupNL reduces training time compared to the vanilla CNN mainly because it reduces the computational workload of the CNN model. This shows that the GroupNL can effectively speed up the training of the model by using nonlinear transformation to generate feature maps to reduce the use of sliding windows to generate feature maps, and also makes the GroupNL-based model more suitable for deployment on IoT devices.
TABLE 11: Training Settings in DDP Evaluation.
Platform Parallelism No. of GPUs LR No. of BS
RTX 2080Ti baseline 1-GPU 0.1 128$\cdot$1
$\times$2 2-GPUs 0.2 128$\cdot$2
$\times$4 2-GPUs 0.4 256$\cdot$2
$\times$8 4-GPUs 0.8 256$\cdot$4
RTX 4090 $\times$8 8-GPUs 0.001 128$\cdot$8
In general, GroupNL can accelerate model training because: (i) the non-trainable parameters contained in GroupNL do not need to be transferred between multiple GPUs, reducing the transmission time during model training; (ii) nonlinear transformation operations replace most operations based on sliding window feature map generation, reducing the calculation time during model training. GroupNL optimizes the model acceleration in terms of parameter transmission and calculation time. Therefore, it can accelerate model training.

### v2

@@ 4.4.3 Module-Level Profiling
TABLE 16: Module-level Profiling on Raspberry Pi with PyTorch-CPU Backend.
Module $r$ #Params #FLOPs (M) Avg.Times (ms) FPS ($\uparrow$) Power (W) Energy (Wh)
Conv - 2.36M 603.98 122.2 8.18 5.003 0.0208
SineFM Conv 2 1.18M 302.97 97.1 10.3 5.396 0.0183
4 0.59M 151.81 55.36 18.06 5.285 0.0112
8 0.30M 76.23 34.86 28.69 5.024 0.0073
Ghost Conv 2 1.18M 302.58 86.31 11.59 5.262 0.0161
4 0.59M 151.88 49.95 20.02 5.156 0.0099
8 0.30M 76.53 31.61 31.63 5.292 0.0072
GroupNL Conv 2 1.18M 302.06 70.14 14.26 5.005 0.0128
4 0.59M 151.09 42.32 23.63 4.905 0.0082
8 0.29M 75.61 28.43 35.17 4.905 0.0063
Depthwise Conv - 5.12K 1.18 48.71 20.53 6.120 0.0116
GroupNL Conv (sparse) 2 4.62K 1.25 30.68 32.6 6.096 0.0081
4 4.63K 1.28 22.65 44.16 5.730 0.0062
8 4.66K 1.29 18.82 53.14 5.831 0.0056
Figure 12: Power in Module-level Profiling.
Figure 13: Energy Consumption in Module-level Profiling.
TABLE 17: Severity Levels Comparison in EfficientNet-ES Structure (Top: Vanilla; Down: GroupNL) on ImageNet-C.
Level mCE ($\downarrow$) Corruption Error ($\downarrow$)
Gauss. Noise Shot Noise Impulse Noise Defocus Blur Glass Blur Motion Blur Zoom Blur Snow Frost Fog Bright. Contrast Elastic Trans. Pixelate JPEG Comp. Speckle Noise Gauss. Blur Spatter Saturate
Mean 55.033 62.052 63.339 65.818 64.430 73.497 58.716 65.290 59.704 54.999 47.713 30.732 45.922 57.111 60.785 44.755 54.082 60.322 41.914 34.451
53.958 60.843 61.825 65.581 62.466 72.968 58.113 63.870 57.496 54.063 47.302 30.174 46.991 55.622 58.412 43.345 53.372 58.514 40.574 33.670
Severity 1 36.260 36.416 36.568 39.068 43.558 46.032 34.822 48.798 40.4 35.408 34.764 27.09 31.504 32.602 36.878 35.164 35.152 33.166 28.818 32.724
35.599 35.996 36.492 40.402 42.060 44.626 33.856 47.136 38.78 34.846 34.686 26.522 31.214 31.780 35.662 34.420 35.50 32.576 27.910 31.910
Severity 2 45.169 43.776 46.122 52.036 51.022 59.184 43.698 58.968 60.264 49.080 38.620 28.162 34.250 53.546 43.180 38.784 39.126 46.816 35.982 35.586
44.314 43.382 45.804 53.748 49.220 58.044 42.622 57.376 57.550 48.358 39.076 27.544 34.332 51.996 41.150 37.810 39.920 45.180 34.616 34.238
Severity 3 53.486 58.540 60.116 61.688 65.410 83.766 58.614 66.520 56.594 59.804 45.486 29.682 38.798 48.584 54.960 41.356 54.754 61.130 42.196 28.240
52.404 57.376 58.850 62.140 63.276 83.656 57.852 65.074 54.288 58.764 45.746 29.134 39.670 46.558 52.248 40.040 54.288 58.622 40.414 27.676
Severity 4 64.426 77.188 81.608 81.218 76.78 87.732 74.308 73.076 66.678 62.122 51.794 32.464 51.208 63.836 77.158 49.130 64.962 73.144 45.842 33.838
62.845 75.034 78.738 78.738 74.538 87.688 74.090 71.694 63.424 60.934 50.998 31.958 52.784 61.382 72.984 47.538 63.078 70.662 44.416 33.374
Severity 5 75.826 94.338 92.28 95.080 85.380 90.772 82.136 79.088 74.586 68.580 67.90 36.262 73.848 86.988 91.750 59.340 76.416 87.356 56.734 41.868
74.628 92.428 89.24 92.878 83.236 90.826 82.146 78.068 73.436 67.412 66.002 35.710 76.956 86.392 90.014 56.918 74.072 85.530 55.514 41.152
4.4.4 Analysis of Corruption Levels
We report the results of five corruption levels for EfficientNet-ES in Table 17. With very few exceptions (e.g., contrast), GroupNL EfficientNet-ES demonstrates superior robustness across all five severity levels of ImageNet-C corruptions. It achieves a lower mCE and outperforms the vanilla model in the vast majority of corruption types at every severity level. This consistent improvement highlights GroupNL’s effectiveness in enhancing model robustness against a wide range of real-world image distortions.